# Large Language Model-based Decision-making for COLREGs and the Control of Autonomous Surface Vehicles

### v2

@@ 1 INTRODUCTION
The integration of LLMs into autonomous systems has seen growing interest across various domains. Chen et al. [2] introduced an LLM-driven framework for multiple-vehicle dispatching, demonstrating the model’s capability to process real-time traffic data and environmental conditions for optimal routing in urban environments. Their work showed how LLMs can effectively handle complex traffic scenarios while maintaining safety constraints. A comprehensive survey [3] explored multimodal LLMs in autonomous driving, highlighting the challenges and opportunities in integrating diverse sensory information for robust decision-making. This integration of multiple data modalities with LLM reasoning capabilities has proven particularly effective in dynamic environments. Building on these foundations, several groundbreaking approaches have emerged. DrPlanner [7] demonstrated the effectiveness of LLMs in diagnosing and repairing motion planners, showing how natural language processing can enhance system reliability and safety. Significant advances in traffic management were achieved through human-mimetic traffic signal control [10], while [11] focused on policy adaptation across diverse driving environments, demonstrating LLMs’ flexibility in handling varying conditions. DILU [12] presented a knowledge-driven approach incorporating human like reasoning, addressing one of the key challenges in autonomous systems, the integration of human-like understanding with machine precision. Another research is LanguageMPC [13], which showed how LLMs can enhance autonomous systems’ ability to interpret complex rules while maintaining safety and efficiency. This work demonstrated the potential for LLMs to bridge the gap between regulations written for human consumption and machine execution. In [14], the authors presented a ‘Drive like a Human’ approach by demonstrating LLMs’ capability to mirror human driver behaviour through natural language reasoning, achieving a level of efficiency in decision-making previously difficult to attain with traditional approaches.
@@ 2 Problem statement: Mission planning and COLREGs-based decision-making
Recent advancements in LLMs have demonstrated remarkable success in complex decision-making tasks, particularly in autonomous driving applications [2, 3, 7, 10, 11, 17] where they enable human-like reasoning and clear explanation of actions. Studies have shown LLMs’ capability to interpret traffic rules, assess dynamic situations, and generate appropriate responses in road environments. While autonomous driving have seen substantial integration of LLMs in their decision-making algorithms , with numerous successful implementations and ongoing research, the application to maritime collision regulations (COLREGs) compliance remains largely unexplored. Also, The absence of robust, human-like decision-making capabilities in ASVs poses significant safety risks as the industry moves toward higher levels of automation. Traditional rule-based systems often struggle with the nuanced interpretation of COLREGs rules and fail to provide clear explanations for their actions, limiting trust and adoption in practical applications. There is a critical need for systems that can combine sophisticated situational awareness with clear, explainable decision-making that mirrors trained mariners expertise.
@@ 3.1 LLM-based decision-making algorithm
The proposed COLREGs-compliant decision-making system employs a novel architecture that leverages LLMs for maritime decision-making. Our system utilised OpenAI’s GPT-4 model accessed through the ChatGPT API, with a temperature setting of 0.2 to ensure consistent and deterministic outputs. The LLM was not specifically trained on maritime data; instead, we leveraged its pre-trained and general reasoning capabilities through careful prompt engineering that explicitly encoded COLREGs rules and maritime navigation requirements. The complete implementation is available online via GitHub111 https://github.com/Klins101/LLM-Based-COLREGs-Decision-Making, including all Python code, prompt templates, and simulation environments used in this study. The architecture consists of four primary components that work in concert to generate safe and explainable navigation decisions as illustrated in Fig. 2. At the highest level, the Task Description and Prompt Engineering component establishes the fundamental framework for the LLM’s operation. This component transforms maritime navigation expertise into a structured prompt that guides the LLM’s understanding of COLREGs rules, relative heading interpretations, and decision-making requirements.
@@
The core of our architecture lies in the LLM-based Decision Maker, which processes current encounter conditions to generate appropriate actions. This component evaluates a specific set of parameters: the relative heading ($\psi_{rel}$) between own vessel and target vessel, risk assessment (Risk) calculated through fuzzy logic, Distance to Closest Point of Approach ($D_{CPA}$), Range, and Time to Closest Point of Approach ($T_{CPA}$). Additionally, it considers the current maneuver status (give-way/stand-on) to maintain decision consistency. By framing these parameters as a condition-based query, the LLM determines the appropriate COLREGs-compliant action for the current situation. The Explainable Action component ensures that all decisions are transparent and interpretable. It generates standardised outputs that include situation classification (overtaking, head-on, or crossing), specific action commands (stand-on or give-way), and clear reasoning chains explaining the decision basis as illustrated in Fig. 5. Supporting these primary components, the maneuver or memory block retains the encounter state and ensures decision consistency throughout maneuvers. It stores crucial information about the initial situation classification, current action status, and navigation history. This state awareness prevents oscillatory behaviour and ensures smooth and predictable decision-making. The Maritime Environment component provides continuous feedback about encounter geometry and risk evolution, enabling the system to adapt to changing conditions while maintaining COLREGs compliance.
@@ 3.1.2 Risk Assessment Integration
$\footnotesize\mathcal{T}=\begin{cases}Risk:0.75,\\ R:1000\ \text{m},\\ D_{CPA}:250\ \text{m}\\ T_{CPA}:60\ \text{s}\end{cases}$ (3)
@@ 3.1.3 State-Aware Decision Maintenance
The system maintains encounter state through a state vector $S$:
@@
  • $i\in\mathbb{N}$: maneuver initiation index
@@ 3.2 The environment and low-level algorithms’ simulation
The environment simulation includes the vehicle model, low-level local planning, control algorithm, and risk assessment, as discussed in [18]. A brief explanation of the process is provided below. The Python code used to simulate the system is available online via GitHub222 https://github.com/Psarhadi/Autonomous_ship_planning_collision_avoidance_control
@@ 4 Simulation Results
In this section, we evaluate the performance of the proposed risk-aware LLM-based algorithm for four typical maritime encounter scenarios. The simulations were conducted at a higher frequency, with the decision-maker operating at 1 Hz in an online environment. In all simulations, the own ship maintained a constant cruising speed of 32 knots (16 m/s), chosen to reflect the operational challenges of a high-speed vessel, with possible collision. The traversed paths for the ASV (black) and the target ship (coloured based on the risk) for all cases are depicted in Fig. 4. In the crossing give-way scenario, the system demonstrated appropriate Rule 16 compliance by initiating a timely starboard turn when collision risk became significant. For head-on encounters, the system exhibited correct interpretation of Rule 14 requirements. When presented with critical parameters indicating an imminent head-on situation, the LLM generated clear decisions for starboard turn manoeuvres, ensuring mutual action compliance as required by COLREGs and as shown in Fig. 4. The overtaking scenario highlighted our system’s capacity for sustained decision consistency. The LLM maintained appropriate Rule 13 compliance throughout the extended manoeuvre, as evidenced by its explicit reasoning about overtaking responsibilities and required actions as specified above.
@@
Notably, In the crossing stand-on situation, the algorithm effectively implemented Rule 17. Despite moderate-risk inputs, the LLM maintained correct stand-on behaviour with the target on the port side, providing clear explanations based on Rule 17. These results were achieved through trial and error and supervising incorrect decisions, but with the proposed training strategy (Fig. 2), the algorithm is able to make online, in-the-loop decisions throughout testing. We further discuss the results of our prompt engineering for two critical situations.
Fig. 5 illustrates a moment in the crossing scenario, demonstrating the LLM’s decision-making process in a critical situation. The trajectory plot (left) shows a developing crossing scenario with converging vessels. The Lalgorithm processes a set of urgent parameters: high risk value of 0.86, critically low $T_{\text{CPA}}$ of 16.84 s, $D_{\text{CPA}}$ of 257.54 m, range 431.79 m, and relative heading of -150.09 degrees. Through the prompt engineering framework (centre), the LLM correctly identifies this as a crossing situation where the own vessel is the give-way vessel. The relative heading of -150.09° (normalised to 209.91°) places the target vessel on our starboard side, making the own vessel the give-way vessel under Rule 15. The Explainable Action(s) panel (right) shows the LLM’s clear reasoning process, resulting in the decision to ‘Give-way, turn starboard’. This action is mandated by both our give-way obligation under Rule 15 and the critical nature of the situation, as evidenced by the high risk value, close range, and critically low $T_{CPA}$. This scenario verifies the system’s ability to correctly interpret relative bearings for proper COLREGs classification while also considering the urgency of the situation in its decision-making process. The explicit explanation referencing the critical parameters demonstrates capability to provide clear reasoning chains for its decisions.
@@
For the crossing stand-on scenario illustrated in Fig. 6, we observe another demonstration of the LLM’s decision-making capability. The trajectory plot (left) shows the evolution of a crossing situation where the target vessel approaches from the port side. The LLM-based decision maker processes a set of parameters indicating a manageable situation: calculated risk value of 0.78, $T_{CPA}$ 53.27s, $D_{CPA}$ 296.21m, range 407.47m, and relative heading 125.08. When this data is processed through our prompt engineering framework (centre), the LLM demonstrates precise understanding of COLREGs Rule 17 requirements for stand-on vessels. The Explainable Action(s) panel (right) shows the system’s clear decision-making logic. With a moderate risk value (0.78) and parameters near but not exceeding critical thresholds, the LLM correctly maintains stand-on behaviour as required when the target vessel is on the port side. The explanation specifically acknowledges ’Risk present but manageable, maintaining stand-on duty as target vessel on port side,’ demonstrating the system’s ability to balance risk awareness with proper COLREGs compliance. This scenario particularly highlights our system’s sophisticated state awareness capabilities. Even with elevated risk values, the LLM maintains appropriate stand-on behaviour while monitoring the situation, precisely as required by Rule 17. This demonstrates the successful integration of risk assessment with COLREGs compliance, ensuring predictable behaviour that other vessels can interpret and respond to appropriately.
Figure 6: decision-making, crossing standon
This diversity of appropriate responses across different encounter types, all generated with clear reasoning chains, validates our LLM’s capability to interpret and respond to various maritime encounters while maintaining regulatory compliance.
@@ 5 CONCLUSION
This paper presented a novel method to maritime COLAV for ASVs using an LLM-based algorithm for COLREGs-compliant decision-making. Our approach utilised a combination of a high-level, risk-aware LLM-based decision maker and rule-based low-level algorithms to perform path tracking with safe manoeuvres in collision encounters. This algorithm generates human-like, explainable interpretations in complex scenarios, proposing appropriate actions to achieve the mission objectives. The devised decision-maker employs crucial parameters, including online risk assessment using CPA calculations, range, and bearing to obstacles, to generate runtime decisions with clear explanations of the reasoning process. Simulation results across four fundamental maritime scenarios involving a semi-realistic ship model with manoeuvring constraints, input saturation, disturbances, and non-holonomic behaviour, demonstrate the effectiveness of our approach. The algorithm’s unique capability lies in its ability to spell out the COLREGs rules, making the decision-making process explainable.
The results lead to several important conclusions. First, LLMs can interpret maritime navigation rules and generate COLREGs-compliant decisions with human-like reasoning. Second, the integration of online risk assessment with LLM decision-making enhances situational awareness and transparency of the decision-making process. Third, our risk-aware approach ensures consistent behaviour throughout encounters, preventing oscillatory decisions while maintaining safety. Advanced versions of these algorithms could even serve as captain-assist features within monitoring tools. The exponential progress in the reasoning capabilities of LLMs will further increase the likelihood of deploying such algorithms in these systems. Future studies could focus on providing an advanced framework to test, refine, analyse, and, importantly, verify the algorithm to further strengthen trust in this solution.

### v3

@@ 1 INTRODUCTION
The integration of LLMs into autonomous systems has seen growing interest across various domains. Chen et al. [2] introduced an LLM-driven framework for multiple-vehicle dispatching, demonstrating the model’s capability to process real-time traffic data and environmental conditions for optimal routing in urban environments. Their work showed how LLMs can effectively handle complex traffic scenarios while maintaining safety constraints. A comprehensive survey [3] explored multimodal LLMs in autonomous driving, highlighting the challenges and opportunities in integrating diverse sensory information for robust decision-making. This integration of multiple data modalities with LLM reasoning capabilities has proven particularly effective in dynamic environments. Building on these foundations, several groundbreaking approaches have emerged. Significant advances in traffic management were achieved through human-mimetic traffic signal control [10], while [11] focused on policy adaptation across diverse driving environments, demonstrating LLMs’ flexibility in handling varying conditions. DILU [12] presented a knowledge-driven approach incorporating human like reasoning, addressing one of the key challenges in autonomous systems, the integration of human-like understanding with machine precision. Another research is LanguageMPC [13], which showed how LLMs can enhance autonomous systems’ ability to interpret complex rules while maintaining safety and efficiency. This work demonstrated the potential for LLMs to bridge the gap between regulations written for human consumption and machine execution. In [14], the authors presented a ‘Drive like a Human’ approach by demonstrating LLMs’ capability to mirror human driver behaviour through natural language reasoning, achieving a level of efficiency in decision-making previously difficult to attain with traditional approaches.
@@ 2 Problem statement: Mission planning and COLREGs-based decision-making
Recent advancements in LLMs have demonstrated remarkable success in complex decision-making tasks, particularly in autonomous driving applications [2, 3, 7, 10, 11, 17] where they enable human-like reasoning and clear explanation of actions. Studies have shown LLMs’ capability to interpret traffic rules, assess dynamic situations, and generate appropriate responses in road environments. While autonomous driving have seen substantial integration of LLMs in their decision-making algorithms , with numerous successful implementations and ongoing research, the application to maritime collision regulations (COLREGs) compliance remains largely unexplored. Also, The absence of human-like decision-making capabilities in ASVs poses significant safety risks as the industry moves toward higher levels of automation. Traditional rule-based systems often struggle with the nuanced interpretation of COLREGs rules and fail to provide clear explanations for their actions, limiting trust and adoption in practical applications. There is a critical need for systems that can combine sophisticated situational awareness with clear, explainable decision-making that mirrors trained mariners expertise.
@@ 3.1 LLM-based decision-making algorithm
The proposed COLREGs-compliant decision-making system employs a novel architecture that leverages LLMs for maritime decision-making. Our system utilised OpenAI’s GPT-4 model accessed through the ChatGPT API, with a temperature setting of 0.2 to ensure consistent and deterministic outputs. The LLM was not specifically trained on maritime data; instead, we leveraged its pre-trained and general reasoning capabilities through careful prompt engineering that explicitly encoded COLREGs rules and maritime navigation requirements. The complete implementation is available online via GitHub111github.com/Klins101/CORALL, including all Python code, prompt templates, and simulation environments used in this study. The architecture consists of four primary components that work in concert to generate safe and explainable navigation decisions as illustrated in Fig. 2. At the highest level, the Task Description and Prompt Engineering component establishes the fundamental framework for the LLM’s operation. This component transforms maritime navigation expertise into a structured prompt that guides the LLM’s understanding of COLREGs rules, relative heading interpretations, and decision-making requirements.
@@
The core of our architecture lies in the LLM-based Decision Maker, which processes current encounter conditions to generate appropriate actions. This component evaluates a specific set of parameters: the relative heading ($\psi_{rel}$) between own vessel and target vessel, risk assessment (Risk) calculated through fuzzy logic, Distance to Closest Point of Approach ($D_{CPA}$), Range, and Time to Closest Point of Approach ($T_{CPA}$). Additionally, it considers the current manoeuvre status (give-way/stand-on) to maintain decision consistency. By framing these parameters as a condition-based query, the LLM determines the appropriate COLREGs-compliant action for the current situation. The Explainable Action component ensures that all decisions are transparent and interpretable. It generates standardised outputs that include situation classification (overtaking, head-on, or crossing), specific action commands (stand-on or give-way), and clear reasoning chains explaining the decision basis as illustrated in Fig. 5. Supporting these primary components, the manoeuvre or memory block retains the encounter state and ensures decision consistency throughout manoeuvres. It stores crucial information about the initial situation classification, current action status, and navigation history. This state of awareness prevents oscillatory behaviour and ensures smooth and predictable decision-making. The Maritime Environment component provides continuous feedback about encounter geometry and risk evolution, enabling the system to adapt to changing conditions while maintaining COLREGs compliance.
@@ 3.1.2 Risk Assessment Integration
$\footnotesize\mathcal{T}=\begin{cases}\mathcal{T}_{Risk}:0.75,\\ \mathcal{T}_{R}:1000\ \text{m},\\ \mathcal{T}_{D_{CPA}}:250\ \text{m}\\ \mathcal{T}_{T_{CPA}}:60\ \text{s}\end{cases}$ (3)
@@ 3.1.3 State-Aware Decision Maintenance
The system maintains the encounter state through a state vector $S$:
@@
  • $i\in\mathbb{N}$: manoeuvre initiation index
@@ 3.2 The environment and low-level algorithms’ simulation
The environment simulation includes the vehicle model, low-level local planning, control algorithm, and risk assessment, as discussed in [18]. A brief explanation of the process is provided below. The Python code used to simulate the system is available online via GitHub222github.com/Psarhadi/Autonomous_ship_planning_control.
@@ 4 Simulation Results
In this section, we evaluate the performance of the proposed LLM-based algorithm based on risk awareness in four typical scenarios of maritime encounters. The simulations were conducted at a high frequency, with the decision-maker operating at 1 Hz in an online environment. It should be noted that considering the dynamics of the ship, lower-level control loops run at 100 Hz. However, a frequency of 1 Hz is considered feasible for higher-level decision-making processes. In all simulations, the own ship maintained a constant cruising speed of 32 knots (16 m/s), chosen to reflect the operational challenges of a high-speed vessel, with possible collision. The traversed paths for the ASV (black) and the target ship (coloured based on the risk) for all cases are depicted in Fig. 4. In the crossing give-way scenario, the system demonstrated appropriate Rule 16 compliance by initiating a timely starboard turn when collision risk became significant. For head-on encounters, the system exhibited correct interpretation of Rule 14 requirements. When presented with critical parameters indicating an imminent head-on situation, the LLM generated clear decisions for starboard turn manoeuvres, ensuring mutual action compliance as required by COLREGs and as shown in Fig. 4. The overtaking scenario highlighted our system’s capacity for sustained decision consistency. The LLM maintained appropriate Rule 13 compliance throughout the extended manoeuvre, as evidenced by its explicit reasoning about overtaking responsibilities and required actions as specified above.
@@
Notably, In the crossing stand-on situation, the algorithm effectively implemented Rule 17. Despite moderate-risk inputs, the LLM maintained correct stand-on behaviour with the target on the port side, providing clear explanations based on Rule 17. With the proposed strategy (Fig. 2), the algorithm is able to make online, in-the-loop decisions throughout testing. We further discuss the results of our prompt engineering for two critical situations.
Fig. 5 illustrates a moment in the crossing scenario, demonstrating the LLM’s decision-making process in a critical situation. The trajectory plot (left) shows a developing crossing scenario with converging vessels. The algorithm processes a set of urgent parameters: high risk value of 0.86, critically low $T_{\text{CPA}}$ of 16.84 s, $D_{\text{CPA}}$ of 257.54 m, range 431.79 m, and relative heading of -150.09 . Through the prompt engineering framework (centre), the LLM correctly identifies this as a crossing situation where the own vessel is the give-way vessel. The relative heading of -150.09 (normalised to 209.91) places the target vessel on our starboard side, making the own vessel the give-way vessel under Rule 15. The Explainable Action(s) panel (right) shows the LLM’s clear reasoning process, resulting in the decision to ‘Give-way, turn starboard’. This action is mandated by both our give-way obligation under Rule 15 and the critical nature of the situation, as evidenced by the high risk value, close range, and critically low $T_{CPA}$. This scenario verifies the system’s ability to correctly interpret relative bearings for proper COLREGs classification while also considering the urgency of the situation in its decision-making process.
@@
For the crossing stand-on scenario illustrated in Fig. 6, we observe another demonstration of the LLM’s decision-making capability. The trajectory plot (left) shows the evolution of a crossing situation where the target vessel approaches from the port side. The LLM-based decision maker processes a set of parameters indicating a manageable situation: calculated risk value of 0.78, $T_{CPA}$ 53.27s, $D_{CPA}$ 296.21m, range 407.47m, and relative heading 125.08. When this data is processed through our prompt engineering framework (centre), the LLM demonstrates precise understanding of COLREGs Rule 17 requirements for stand-on vessels. The Explainable Action(s) panel (right) shows the system’s clear decision-making logic. With a moderate risk value (0.78) and parameters near but not exceeding critical thresholds, the LLM correctly maintains stand-on behaviour as required when the target vessel is on the port side. The explanation specifically acknowledges ’Risk present but manageable, maintaining stand-on duty as target vessel on the port side,’ demonstrating the system’s ability to balance risk awareness with proper COLREGs compliance. This scenario particularly highlights our system’s sophisticated state awareness capabilities. Even with elevated risk values, the LLM maintains appropriate stand-on behaviour while monitoring the situation, precisely as required by Rule 17.
Figure 6: decision-making, crossing stand-on
This diversity of appropriate responses across different encounter types, all generated with clear reasoning chains, verifies our LLM’s capability to interpret and respond to various maritime encounters while maintaining regulatory compliance.
@@ 5 CONCLUSION
This paper presented a novel method to maritime COLAV for ASVs using an LLM-based algorithm for COLREGs-compliant decision-making. Our approach utilised a combination of a high-level, risk-aware LLM-based decision maker and rule-based low-level algorithms to perform path tracking with safe manoeuvres in collision encounters. This algorithm generates human-like, explainable interpretations in complex scenarios, proposing appropriate actions to achieve the mission objectives. The devised decision-maker employs crucial parameters, including online risk assessment using CPA calculations, range, and bearing to obstacles, to generate runtime decisions with clear explanations of the reasoning process. Simulation results across four fundamental maritime scenarios involving a semi-realistic ship model with manoeuvring constraints, input saturation, disturbances, and non-holonomic behaviour, demonstrate the effectiveness of our approach.
The unique capability of the algorithm lies in its ability to spell out the COLREG rules, making the decision-making process explainable. The results lead to several important conclusions. First, LLMs can interpret maritime navigation rules and generate COLREGs-compliant decisions with human-like reasoning. Second, the integration of online risk assessment with LLM decision-making enhances situational awareness and transparency of the decision-making process. Third, our risk-aware approach ensures consistent behaviour throughout encounters, preventing oscillatory decisions while maintaining safety. Advanced versions of these algorithms could even serve as captain-assist features within monitoring tools. The exponential progress in the reasoning capabilities of LLMs will further increase the likelihood of deploying such algorithms in these systems. While our fuzzy logic risk assessment model effectively combines navigation parameters into a unified risk metric, but has limitations. It may not capture all collision risk factors such as weather, vessel manoeuvreability, and traffic density, while they could be incorporated into the thresholds as varying parameters.
Future studies should focus on testing and verifying the algorithm to build greater trust in this solution. The model currently lacks consideration of important factors like weather, vessel manoeuvrability and traffic density, which presents opportunities for enhancement in subsequent research.